This Question: 1 pt 3 of 10 (0 This Quiz: 10 pts possible Question Help A random sample of size n 80 yielded p o.60. ct a 90% conf dence interval for p. interval. d Explain what is meant by the phrase 90% confidence interval. a. Is the sample large enough? O A. No, because np c 15 and na 2 15. O B. No, because np <15 and ng < 15 O C. No, because np - 15 and ng < 15 d) D. Yes, because np215and nq215. mpts t of 2 DD b, The 90% confidence interval for p is ( Round to two decimal places as needed.) c. Interpret this confidence interval 0 A. We are confident that 90% of the population is outside re interval for p 0 B. We are confident that 90% of the population is described by the interval forp. ○ C. We are 90% confident that p lies in the confidence interval. 0 D. There is a 90% chance that the value of p is outside the interval d Explain what is meant by the phrase 90% confidence interval- 0 of 6
O C. No, because np $15 and na <15. 0 D. Yes, because np215 and nqz 15. b. The 90% confidence interval for p is ( Round to two decimal places as needed.) c. Interpret this confidence interval. O A. We are confident that 90% of the population is outside the interval for p. B. We are confident that 90% of the population is described by the interval for p. ° C. We are 90% confident that p lies in the confidence interval. D. There is a 90% chance that the value of p is outside the interval. d. Explain what is meant by the phrase 90% confidence interval. A. 0 B. The confidence interval is within 10% of the true parameter v C. D. 90% of similarty constructed intervals would contain the true value of the parameter. 90% of similarly constructed intervals would be centered on the true value of the parameter Tere is a 90% chance that the true value of the parameter is the midpoint of the interval. Click to select your answer(s). 2 6
Answer #1

(a)

n = 80

hat{p} = 0.6

So,

hat{q} = 1 - 0.6 =0.4

So,

np 80 × 0.6 48

ng 80 × 0.4-32

So,

Correct option:

D. Yes, because nnp > 15 and na > 15

(b)

SE = pi/n-/0.6 x 0.4/80 0.0548

alpha = 0.10

From Table, critical values of Z = pm 1.645

Confidence interval:

0.6 pm (1.645 X 0.0548)

= 0.6 pm 0.0901

= (0.51, 0.69)

So,

90% confidence interval for p is:

(0.51,0.69)

(c)

Correct option:

C. We are 90% confident that p lies in the confidence interval.

(d)

Correct option:

B. 90% of similarly constructed confidence intervals would contain the true parameter value.